(1) find the exact value of sec(225degrees)


(2) In the triangle ABC, find A, if B= 100 degrees, and a=20.

(3) In the triangle ABC, find b, iF A=15, C=12 and angle b=150 degrees.

Answer by venugopalramana(3286)

(1) find the exact value of sec(225degrees)=1/COS(225)=1/COS(180+45)=-1/COS(45)
=-SQRT2=-1.414
(2) In the triangle ABC, find A, if B= 100 degrees, and a=20.
DATA IS INSUFFICIENT.CHECK PROBLEM
(3) In the triangle ABC, find b, iF A=15, C=12 and angle b=150 degrees.
b^2=a^+c^2-2acCOS(B)=15^2+12^2-2*15*12COS(150)=225+144-30*COS(180-30)
=369+30COS(30)=369+30*0.866=369+25.98=394.98
b=SQRT(394.98)=19.87